Suppose that poutine ice cream is a hot commodity with the inverse market demand function ?(?) = 50 − ?/100 where ? denotes quarts per week. There is only one confectioner with the secret recipe. Despite their tireless efforts, other confectioners cannot replicate this tasty treat.

A. What kind of market is this?

B. What is the confectioner’s total revenue function?

C. From the total revenue function, we know that marginal revenue is ??(?) = 50 − ?/50 . Moreover, the confectioner has the cost function is ?(?) = 4?. This implies that the marginal cost is $4 per quart. What is the optimal level of output for the confectioner?

D. Under uniform pricing, how much will the confectioner charge per quart? Hint: This ice cream isn’t cheap!

E. The confectioner can earn positive economic profit, even in the long run, because of barriers to entry. What is the confectioner’s profit?

F. With price on the vertical axis and output on the horizontal axis, sketch the demand, marginal revenue and marginal cost curves, as well as the confectioner’s optimal level of output and price. Clearly label the axes, curves, optimal level of output and price.

G. On your graph, show the market output and price under pure competition/perfect competition (e.g. if the recipe became widely known and used by other confectioners). Relative to this outcome, indicate the loss in total surplus, known as deadweight loss, that results from keeping the recipe secret and impossible to replicate.

H. Instead of uniform pricing, suppose the confectioner wants to try to increase profit via price discrimination. They’ve hired you as a consultant. In three sentences of less, suggest one way this confectioner could engage in price discrimination. It can be first, second or third degree price discrimination.
Answer #1

ANSWER;

Suppose that poutine ice cream is a hot commodity with the inverse market demand function

P = 50 - (y/100)

(A) What kind of market is this

Since only one seller exists, this is a monopoly.

(B)What is the confectioner s total revenue function

TR = P x y = 50y - (y2/100)

(C)From the total revenue function

Output is optimal when MR = MC.

50 - (y/50) = 4

y/50 = 46

y = 2300

(D)Under uniform pricing how much will the confectioner  

P = 50 - (2300/100) = 50 - 23 = 27

(E)The confectioner can earn positive economic profit

Profit = y x (P - MC) = 2300 x (27 - 4) = 2300 x 23 = 52900

NOTE: As per Chegg Answering Policy, 1st 5 parts are answered.